OpenStudy (tennistar):

Rewrite f(x) = 3(x − 2)2 + 1 from vertex form to standard form.

OpenStudy (tennistar):

f(x) = 3x2 + 13 f(x) = 3x2 − 12x + 13 f(x) = 3x2 + 12x − 11 f(x) = 9x2 − 36x + 37

OpenStudy (tkhunny):

Have you considered expanding the parentheses and giving it a look?

OpenStudy (tennistar):

Hmm. Yes, but which do I do first? Do I square the entire equation within the parenthesis, or do I distribute?

OpenStudy (xmissalycatx):

Use PEMDAS! You use the order PEMDAS a lot in math. It means Parentheses, Exponent, Multiply, Divide, Add, Subtract.